# LEO-Mini: An Efficient Multimodal Large Language Model using Conditional Token Reduction and Mixture of Multi-Modal Experts

### v1

@@ 4.1 Implementations and Benchmarks
Training. LEO-Mini uses the same training data as LLaVA-v1.5 [39] in stage 3 with 665K instruction data. For stage 1 and stage 2, we follow the same data as EAGLE [54]. The training was conducted on 8 A6000 GPUs (48 GB) using DeepSpeed’s Zero2 strategy [51].
4.2 Main Results
We compare our LEO-Mini with several state-of-the-art token reduction MLLMs (Tab. 3), MoE-based MLLMs (Tab. 3), and general MLLMs (Tab. 3) on various tasks.
Comparison with token reduction MLLMs [20, 52, 33, 63, 31, 68]. The results are shown in Tab. 3. LEO-Mini outperforms all the token reduction MLLMs on all tasks. Specifically, LEO-Mini with Llama3-8B achieves 1583.0 on MME (perception), 77.0 on MMBench, 75.8 on SeedBench, 64.5 on GQA, 69.3 on VizWiz, 75.1 on TextVQA, 84.5 on SQA, 90.3 on POPE, and 38.8 on MMMU, outperforming the best baseline by a large margin. This demonstrates the effectiveness of LEO-Mini in improving the performance of MLLMs on various tasks. Moreover, LEO-Mini with Vicuna-7B also outperforms the best baseline on all tasks except for GQA and MMMU, showing the generalization ability of LEO-Mini with different LLMs.
Model MMEP SEEDI GQA SQA MMMU POPE AI2D TextVQA ChartQA OCRBench
LEO-Mini (64 tokens) 1583 75.8 64.5 84.5 38.8 90.3 75.7 75.1 80.5 62.4
w/ 1 visual token 1565 74.6 64.2 83.6 37.7 89.0 74.5 73.5 80.0 61.7
w/ 16 visual token 1577 75.4 64.4 84.3 38.2 90.1 75.4 74.2 80.2 62.6
w/ 256 visual token 1584 76.1 64.2 85.5 39.0 90.7 75.5 75.5 80.5 63.0
w/ 1.8m SFT data (stage 3) 1548 76.6 64.1 85.9 39.4 90.3 77.9 75.7 80.9 63.1
Comparison with general and MoVE-based MLLMs [8, 39, 40, 5, 37, 35, 47, 34, 11, 1, 45, 57, 54]. The results are shown in Tab. 3. With Vicuna-7B, LEO-Mini outperforms all the general and MoVE-based MLLMs on MME, POPE, AI2D, and OCRBench by 15, 1.2, 2.6, and 2.7 points, respectively, with only 64 visual tokens. LEO-Mini with Vicuna-7B also achieves the second best performance on SeedBenchI, SQA, and TextVQA. On the other side, with Llama3-8B, LEO-Mini achieves stronger performance as it outperforms all the general and MoVE-based MLLMs on MMBench, SQA, and CharQA by 1.1, 0.2, and 0.4 points, respectively, while reducing the number of visual tokens by 97.77% compared to LLaVA-NeXt and Mini-Gemeni-HD. Moreover, compared to the general MLLMs, LEO-Mini with Llama3-8B achieves comparable performance on other benchmark datasets as the discrepancy is marginal. Specifically, LEO-Mini with Llama3-8B achieves the second-best performance on SeedBenchI, AI2D, TextVQA, DocVQA, and OCRBench.
Figure 5: Visualization of the expert choice using LEO-Mini-Llama-8B on TextVQA, ScienceQA, and GQA. Best viewed in color.
4.3 Ablation Studies
In this section, we conduct ablation studies to analyze the effectiveness of different components in LEO-Mini, the number of visual tokens, and the amount of training data for stage 3 using LEO-Mini-Llama-8B. Ablations using other token reduction modules and MoE are presented in the Appendix.
Is 1 visual token enough for MLLMs? To understand whether 1 visual token is enough for representing the visual information, we conduct the experiments and present the results in Tab. 4. LEO-Mini with 1 visual token ($N^{V}=1$) shows a slight decrease across some metrics compared to LEO-Mini with 64 tokens, which is reasonable. For example, in the MMEP, the score dropped from 1583 to 1565, and in SEEDI, the performance slightly decreased from 75.8 to 75.6. Similarly, in the domains of GQA and SQA, the scores decrease from 64.5 to 64.2 and 84.5 to 83.6, respectively. This trend continues across other evaluated areas such as MMMU, POPE, AI2D, TextVQA, ChartQA, and OCRBench. While the use of only 1 visual token consistently leads to lower performance, it greatly improves model’s efficiency as shown in Tab. 5. To further understand how increasing the number of visual tokens impacts the performance, we conduct experiments with 16 and 256 visual tokens. The results show that the performance on most of the benchmarks is improved as the number of visual tokens increases. This is reasonable as more visual tokens can provide more detailed visual information to the model, which can help the model to better understand the visual information. However, we also observe that the performance on some benchmarks is slightly decreased, e.g., AI2D, which might be due to the overfitting issue.
[…]
Models # of VT FLOPs (T) CUDA Time (s)
LLaVA-Next [40] 2880 45.5 7.037
Eagle-X4-Plus [54] 1024 29.4 2.073
LEO-Mini-Llama-8B 1 13.7 0.593
LEO-Mini-Llama-8B 64 16.4 0.655
$\Delta$downarrow 97.77% $\downarrow$ 63.83% $\downarrow$ 90.69%
Figure 6: Qualitative results across four vision-language tasks demonstrating LEO-MINI’s detailed visual understanding. The images are taken from TextVQA [55], DocVQA [49], and MMMU [64].
How does MMoE-LLM switch between different LORA experts? To understand how our MMoE-LLM switches between different LoRA experts, we visualize the expert choice in Fig. 5 (full results are presented in the Appendix due to space constraints) with GQA (general), TextVQA (OCR), and SQA (Knowledge). We observe that the model effectively switches between different LoRA experts based on the input data. For example, in the TextVQA and SQA tasks, the model mainly activates expert 2 at layer 0, while for GQA, the model evenly activates three experts. When it comes to the middle layer, such as layer 29, the model evenly activates experts 0 and 1 for TextVQA and GQA, while for SQA, the model mainly activates expert 1, respectively. This indicates that MMoE-LLM effectively utilizes the multi-modal input instructions to switch between different experts.
4.4 Efficiency Analysis
To understand how efficient LEO-Mini is, we compare the number of visual tokens, FLOPs, and CUDA processing time of LEO-Mini with other MLLMs [40, 54, 57] using Llama3-8B. Specifically, we use Pytorch Profiler to measure the FLOPs and CUDA time using the figure and “What is shown in this image?” as input. We use one A6000 GPU with 48 GB memory to inference the models.
The comparison is shown in Tab. 5. LEO-Mini with Llama3-8B is 97.77% more efficient in terms of visual tokens, 63.83% in terms of FLOPs, and 90.69% in terms of CUDA time compared to LLaVA-Next. Moreover, as shown in Fig. 1, LEO-Mini outperforms LLaVA-Next by 6% on average performance with only 64 visual tokens. Even compared to the most powerful MLLM with similar performance, i.e., Eagle-X4-Plus, LEO-Mini is 93.75% more efficient in terms of visual tokens, 44.10% in terms of FLOPs, and 68.40% in terms of CUDA time. More excitingly and importantly, LEO-Mini with only 1 token achieves comparable performance to the state-of-the-art MLLMs with thousands of visual tokens with faster inference time and lower computational cost in terms of FLOPs.
4.5 Qualitative Analysis
To understand the detailed visual understanding of LEO-Mini, we conduct a case study for qualitative analysis on four vision-language tasks [55, 49, 64] as shown in Fig. 6. We use LEO-Mini-Llama3-8B with 64 tokens.
Overall, LEO-Mini shows strong visual understanding with improved efficiency, making it a practical solution for multi-modal understanding.
5 Conclusion and Limitations
In this paper, to address the redundancy of visual tokens in MLLMs, we propose a novel MLLM, LEO-Mini, that significantly reduces the number of visual tokens while boosting visual reasoning capabilities. LEO-Mini incorporates a novel token reduction module, CoTR, to consolidate a large number of visual tokens into a smaller set of tokens, using the similarity between visual tokens, text tokens, and a compact learnable query. However, simply reducing the number of visual tokens leads to an information loss. To avoid the loss and boost the visual comprehension ability with minimal computational overhead, LEO-Mini employs a novel mixture of multi-modal experts module, MMoE, that includes a set of language (LoRA) experts and a set of vision experts trained on diverse domain-specific data. For better switching between different LoRA experts, MMoE employs a new router that takes the text and visual tokens as additional inputs. MMoE also includes a general LoRA expert that is always activated to learn general knowledge. We evaluate LEO-Mini on various vision-and-language tasks, showcasing its potential for practical applications with improved efficiency and performance compared to existing efficient MLLMs.
Limitations. First, the proposed CoTR needs training. Introducing a training-free token reduction module might be a promising direction. Second, the proposed MMoE is designed to be efficient and scalable, but it may not be optimal for all tasks. For vision experts, due to the computational limitations, we use a fixed set of experts. It would be interesting to explore more vision experts from a wide range of domains and introduce more advanced expert selection mechanisms, such as dynamic vision expert routing. Last, due to the computational limitations, we did not test the efficiency of LEO-Mini using bigger LLMs with 13B or 67B parameters.
Supplementary Material
Token Reduction MoE MMEP SEEDI GQA SQA MMMU POPE AI2D TextVQA ChartQA OCRBench
CoTR MMoE-Vision + MMoE-LLM 1583 75.8 64.5 84.5 38.8 90.3 75.7 75.1 80.5 62.4
MQT MMoE-Vision + MMoE-LLM 1537 75.3 64.4 84.2 37.0 90.4 75.3 74.2 80.4 61.6
MQT - 1435 - 61.6 67.6 34.8 84.4 - - - -
CoTR MMoE-Vision + LoRA-MoE 1521 75.3 64.5 83.4 38.2 90.3 75.4 74.0 80.0 61.7
Table 6: Ablation study on the effectiveness of CoTR and MMoE-LLM. We use Llama3-8B as the backbone.
Figure 7: Visualization of the expert choice using LEO-Mini-Llama-8B on TextVQA, ScienceQA, and GQA of layers 0 to 23. Best viewed in color.
Figure 8: Visualization of the expert choice using LEO-Mini-Llama-8B on TextVQA, ScienceQA, and GQA of layers 24 to 31. Best viewed in color.
Appendix A Experiments
A.1 Benchmark Datasets
We evaluate our method on the following benchmark datasets: MME [14], MMBench [41], Seed-Bench [29], GQA [22], SQA [46], MMMU [64], POPE [32], AI2D [26], VizWiz [16], TextVQA [55], DocVQA [49], ChartQA [48], and OCRBench [42].
MME [14]. The MME benchmark is designed to rigorously evaluate a model’s perceptual and cognitive abilities through 14 subtasks. It employs carefully constructed instruction-answer pairs and concise instructions to minimize data leakage and ensure fair evaluation. This setup provides a robust measure of a model’s performance across various tasks.
MMBench [41]. MMBench offers a hierarchical evaluation framework, categorizing model capabilities into three levels. The first level (L-1) focuses on perception and reasoning. The second level (L-2) expands this to six sub-abilities, while the third level (L-3) further refines these into 20 specific dimensions. This structured approach allows for a nuanced and comprehensive assessment of a model’s multifaceted abilities.
Seed-Bench [29]. SEED-Bench consists of 19K multiple-choice questions with accurate human annotations, covering 12 evaluation dimensions including both the spatial and temporal understanding.
GQA [22]. GQA is structured around three core components: scene graphs, questions, and images. It includes not only the images themselves but also detailed spatial features and object-level attributes. The questions are crafted to assess a model’s ability to comprehend visual scenes and perform reasoning tasks based on the image content.
ScienceQA [46]. ScienceQA spans a wide array of domains, including natural, language, and social sciences. Questions are hierarchically categorized into 26 topics, 127 categories, and 379 skills, providing a diverse and comprehensive testbed for evaluating multimodal understanding, multi-step reasoning, and interpretability.
MMMU [64]. MMMU includes 11.5K meticulously collected multimodal questions from college exams, quizzes, and textbooks, covering six core disciplines: Art & Design, Business, Science, Health & Medicine, Humanities & Social Science, and Tech & Engineering. These questions span 30 subjects and 183 subfields, comprising 30 highly heterogeneous image types, such as charts, diagrams, maps, tables, music sheets, and chemical structures.
POPE [32]. POPE is tailored to assess object hallucination in models. It presents a series of binary questions about the presence of objects in images, using accuracy, recall, precision, and F1 score as metrics. This approach offers a precise evaluation of hallucination levels under different sampling strategies.
AI2D [26]. AI2D is a dataset of over 5000 grade school science diagrams with over 150000 rich annotations, their ground truth syntactic parses, and more than 15000 corresponding multiple choice questions.
VizWiz [16]. VizWiz consists of over 31,000 visual questions originating from blind people who each took a picture using a mobile phone and recorded a spoken question about it, together with 10 crowdsourced answers per visual question.
TextVQA [55]. TextVQA emphasizes the integration of textual information within images. It evaluates a model’s proficiency in reading and reasoning about text embedded in visual content, requiring both visual and textual comprehension to answer questions accurately.
DocVQA [49]. DocVQA consists of 50,000 questions defined on 12,000+ document images.
ChartQA [48]. CharQA is a large benchmark covering 9.6K human-written questions as well as 23.1K questions generated from human-written chart summaries.
OCRBench [42]. OCRBench is a comprehensive benchmark for evaluating the OCR capabilities of multi-modal language models across five key tasks: text recognition, scene text-centric and document-oriented VQA, key information extraction, and handwritten mathematical expression recognition.
A.2 Ablation Studies
@@
How do other token reduction methods work? To understand how our proposed token reduction module, i.e., CoTR, works, we compare it with a representative token reduction method, i.e., MQT-LLaVA [20], as shown in Tab. 6, while keeping the MMoE unchanged. The results show that CoTR outperforms MQT on most benchmarks, demonstrating the effectiveness of our proposed method. Moreover, compared to MQT-LLaVA (the third row), MMoE improves the performance of MQT-LLaVA on most benchmarks, showing the importance of introducing multiple experts for extracting informative tokens.
@@
How does MMoE-LLM improve the performance? We conduct an ablation study to investigate the effectiveness of MMoE-LLM in Tab. 6. Results show that MMoE-LLM significantly improves the performance on most benchmarks, compared to the baseline with MMoE-Vision and LoRA-MoE [61]. MMoE-LLM improves performance by 62, 0.5, 1.1, 0.6, 0.3, 1.1, 0.5, and 0.7 on MMEP, SEEDBenchI, SQA, MMMU, AI2D, TextVQA, ChartQA, and OCRBench, respectively. This demonstrates the effectiveness of MMoE-LLM in better switching between different experts to better understand visual information.
@@
How does MMoE-LLM switch between different LORA experts (full results)? The full results are presented in Figs. 7 and 8. The visualization shows that MMoE-LLM can effectively switch between different experts to better understand visual information. We notice that for some layers, such as layers 0, 3, 5, 6, 23, 25, 29, and 31, MMoE-LLM selects different experts for different benchmarks, while for some layers, such as layers 1, 2, 17, and 27, the model selects the similar experts for different benchmarks. This demonstrates the effectiveness of MMoE-LLM in better understanding visual information by switching between different experts.

### v2

@@ A.2 Training Details
LEO-Mini uses the same training data as LLaVA-v1.5 Liu et al. (2024b) in stage 3 with 665K instruction data. For stage 1 and stage 2, we follow the same data as EAGLE Shi et al. (2024). The training was conducted on 8 A6000 GPUs (48 GB) using DeepSpeed’s Zero2 strategy Rajbhandari et al. (2020).
A.3 Ablation Studies
@@
How do other token reduction methods work? To understand how our proposed token reduction module, i.e., CoTR, works, we compare it with a representative token reduction method, i.e., MQT-LLaVA Hu et al. (2024), as shown in Table˜5, while keeping the MMoE unchanged. The results show that CoTR outperforms MQT on most benchmarks, demonstrating the effectiveness of our proposed method. Moreover, compared to MQT-LLaVA (the third row), MMoE improves the performance of MQT-LLaVA on most benchmarks, showing the importance of introducing multiple experts for extracting informative tokens.
@@
How does MMoE-LLM improve the performance? We conduct an ablation study to investigate the effectiveness of MMoE-LLM in Table˜5. Results show that MMoE-LLM significantly improves the performance on most benchmarks, compared to the baseline with MMoE-Vision and LoRA-MoE Wu et al. (2023). MMoE-LLM improves performance by 62, 0.5, 1.1, 0.6, 0.3, 1.1, 0.5, and 0.7 on MMEP, SEEDBenchI, SQA, MMMU, AI2D, TextVQA, ChartQA, and OCRBench, respectively. This demonstrates the effectiveness of MMoE-LLM in better switching between different experts to better understand visual information.
@@
How does MMoE-LLM switch between different LORA experts (full results)? The full results are presented in Figure˜7. The visualization shows that MMoE-LLM can effectively switch between different experts to better understand visual information. We notice that for some layers, such as layers 0, 3, 5, 6, 23, 25, 29, and 31, MMoE-LLM selects different experts for different benchmarks, while for some layers, such as layers 1, 2, 17, and 27, the model selects the similar experts for different benchmarks. This demonstrates the effectiveness of MMoE-LLM in better understanding visual information by switching between different experts.
How does MMoE-LLM switch between different LORA experts compared to the vanilla MoE? To understand how MMoE-LLM switches using the additional visual and input signals, we compare it with the vanilla MoE in Figure˜8. The full routing results of the vanilla MoE are shown in Figure˜9, while the results of MMoE-LLM are shown in Figure˜7. The results show that at different layers, MMoE-LLM shows different experct preferences for different benchmarks, while the vanilla MoE shows similar expert preferences for different benchmarks.